OpenStudy (anonymous):

What are the factors of 3x^2 + 10x + 3? (3x - 1)(x - 3) (3x - 3)(x - 1) (3x + 1)(x + 3) (3x + 3)(x + 1)

OpenStudy (anonymous):

The left hand side factors into a product with two terms: (x+3) (3 x+1) = 0 Solve each term in the product separately. Split into two equations: x+3 = 0 or 3 x+1 = 0 Look at the first equation: Solve for x. Subtract 3 from both sides: x = -3 or 3 x+1 = 0 Look at the second equation: Isolate terms with x to the left hand side. Subtract 1 from both sides: x = -3 or 3 x = -1 Solve for x. Divide both sides by 3: x = -3 or x = -1/3

OpenStudy (anonymous):

\(\large 3x^2+10x+3\\\large =3x^2+9x+x+3\\\large =3x(x+3)+(x+3)\\\large =(x+3)(3x+1)\)

OpenStudy (anonymous):

Thank you @JungHyunRan and @gowtham89 !

OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 49x^2 - 28x + 16 4a^2 - 20a + 25 25b^2 - 20b - 16 16x^2 - 24x - 9

OpenStudy (anonymous):

answer is 4a^2 - 20a + 25 since the Delta = 0

OpenStudy (anonymous):

I figured that one out right as you sent it. Haha.

OpenStudy (anonymous):

Which of the following is the graph of f(x) = x^2 + 3x - 4? graph of a quadratic function with a minimum at 2, negative 9 and x intercepts at negative 1 and 5 graph of a quadratic function with a minimum at 3, negative 4 and x intercepts at 1 and 5 graph of a quadratic function with a minimum at 2.5, negative 2.4 and x intercepts at 1 and 4 graph of a quadratic function with a minimum at negative 1.5, negative 6.2 and x intercepts at 1 and negative 4

OpenStudy (anonymous):

graph of a quadratic function with a minimum at negative 1.5, negative 6.2 and x intercepts at 1 and negative 4

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

u r welcome buddy